# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

### 16 Mar 2026 v1

@@
Abstract. A deep backward regression-based (DBR) scheme for solving high-dimensional nonlinear parabolic partial differential equations is proposed. Building upon the established DBDP method (Huré et al., 2020), our algorithm introduces a reformulation of the local loss functions that are sequentially optimized via backward induction at each time step. The core of this approach involves reformulating simulated backward stochastic difference equations into their conditional expectation representations, thereby recasting a projection-based stochastic optimization problem as a deterministic function-approximation task. By explicitly incorporating conditional expectations, the DBR scheme facilitates a denoising mechanism prior to loss evaluation. This architecture substantially mitigates numerical variance, resulting in enhanced training stability and superior generalization performance. Numerical results demonstrate that the DBR scheme consistently outperforms the DBDP1 method, maintaining accuracy up to $d=200$ for bounded solutions (see Table 1). Notably, for complex unbounded PDEs where the DBDP1 method fails beyond $d=10$, the DBR scheme remains robust up to $d=20$ with relative errors under $9.7\%$ (see Table 6). Theoretically, we derive rigorous upper error bounds and establish half-order convergence for the proposed scheme. Extensions to variational inequalities are also provided.
Key words. high dimensional nonlinear PDEs, backward stochastic differential equations, conditional expectations, deep neural networks, numerical approximations
AMS subject classifications. 60H35, 65C20, 65M12
@@ 1 Introduction
In this paper, we develop a class of artificial neural networks based on probabilistic numerical methods to approximate the solutions of high-dimensional quasilinear parabolic PDEs of the following form
@@
The pioneer works of Pardoux and Peng [31, 32], Peng [33] show that the quasilinear parabolic PDEs (1.1) are associated with Markovian forward backward stochastic differential equations (FBSDEs) due to the nonlinear Feynman-Kac formula. Since then, significant interest has been directed toward the probabilistic interpretation of PDE solutions. By these interpretations, researchers have developed probabilistic numerical methods to approximate PDE solutions (see [9, 10, 12] and references therein). Concurrently, the design of algorithms to solve FBSDEs has been intensively studied. Up to now, the proposed numerical algorithms to solve nonlinear FBSDEs have shown their remarkable performance with respect to the high-order problems. For instance, the Euler schemes with convergence rates of $\frac{1}{2}$ or $1$ are presented in [2, 19, 36, 37]. Multi-step schemes [24, 25, 21, 38] achieve convergence rates greater than or equal to $1$, although they require terminal conditions over multiple steps. Furthermore, one-step schemes with convergence rates of at least $2$ have been proposed in [8, 35, 26].
@@
However, most of these numerical methods are unsuitable for solving nonlinear PDEs in dimensions greater than 4 (see [4]). To address this issue, some techniques such as the parallel computing method, the sparse grid method, and the variance reduction method(including the control variate methods, the importance sampling methods, and the multilevel Monte Carlo methods et al.) have proven effective for nonlinear PDEs with the number of dimension below 20 (see [20]). The high dimensional problem (that is the so-called curse of dimensionality) remains a major challenge in the numerical solutions of PDEs, even in the broader field of scientific computing.
@@
Until 2017, E, Han and Jentzen [11, 22] introduce the first deep learning based on numerical method, called the deep BSDE method which uses the required values as parameters and utilizes the feedforward neural networks to approximate the solutions through solving an optimization problem and demonstrate the effectiveness in 100-dimensional settings by conducting various numerical experiments. To date, there are lots of numerical algorithms based on neural networks. They are designed by adjusting, reformulating or extending the deep BSDE method to solve high dimensional PDEs (see [13, 17, 28, 23, 29, 34, 15, 16] and many others) and these deep numerical algorithms have also shown their remarkable performance on the high-dimensional problems. Recently, Cai, Fang and Zhou [5, 7] introduce the SOC-MartNet method to solve ultra-high-dimensional quasilinear parabolic PDEs and it demonstrates the strong performance on large-scale benchmarks. Furthermore, they propose the deep random difference method in [6] to reduce variance and improve stability of the SOC-MartNet method.
@@
Among the literature on deep learning algorithms for high-dimensional PDEs, an important DBDP method was proposed by Hure, Pham and Warin in [28]. It estimates the solutions and their gradient by the loss functions, for $i=N-1,N-2,\cdots,0$
$J_{i}^{HPW}(\theta_{i})=\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{Y}_{i+1}^{*}-hf(t_{i},X_{i}^{\pi},\mathcal{Y}_{i},\mathcal{Z}_{i})+\mathcal{Z}_{i}\Delta W_{i}\right|^{2}\right],$ (1.2)
@@
at each time step $t_{i}$, where $\mathcal{Y}_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ and $\mathcal{Z}_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n\times d}$ denote the sequences of network functions. Then update $(\mathcal{Y}_{i}^{*},\mathcal{Z}_{i}^{*})$ as the solutions of the local minizamition problems. Thus, we only need to consider the local gradient as updating the parameter of neural networks, which significantly reduces the difficulty of backpropagation and solves the training difficulties brought by long-sequence time dependencies. The loss function (1.2) is based on the Euler discretization of the BSDE. It measures the residual of the discretized Euler scheme which implies that during the training process, one can directly assess the accuracy of the scheme by observing whether the loss value approaches zero. But the random noise term $\mathcal{Z}_{i}\Delta W_{i}$ in (1.2) may lead to fluctuations especially when $\mathcal{Z}_{i}$ is large (such as high volatility or complex hedging requirements). Naturally, it may cause oscillations in the gradients which lead to divergence. Besides, the loss function $J_{i}^{HPW}(\theta_{i})$ is prone to overfitting the noise of the simulated paths, leading to poor generalization ability.
@@
$F_{y,i}(\theta_{y,i})=\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}^{*}+hf(t_{i},X_{i}^{\pi},\mathcal{Y}_{i},\mathcal{Z}_{i}^{*})\right]\right|^{2}\right],$ (1.3)
@@
$F_{z,i}(\theta_{z,i})=\mathbb{E}\left[\left|\mathcal{Z}_{i}-\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}^{*}\frac{\Delta W_{i}^{\top}}{h}\right]\right|^{2}\right],$ (1.4)
@@
where $\mathbb{E}_{i}\left[\cdot\right]=\mathbb{E}\left[\cdot\big|\mathcal{F}_{t_{i}}\right]$ is the conditional expectation. The above regression-based loss functions which utilize conditional expectations are different from those proposed in Remark 2.3 of Germain, Pham and Warin [17]. The accuracy of the DBR scheme can also be tested when computing at each time step the infimum of loss function, which should be equal to zero as the DBDP scheme in [17]. By explicitly incorporating the conditional expectations, the DBR scheme effectively performs a ‘denoising’ step prior to loss evaluation. In terms of implementation, this is typically realized through multi-path Monte Carlo averaging. Its essence lies in replacing the simulated backward stochastic difference equations with their conditional expectation forms and transforming a projection-based stochastic optimization problem into a smoother deterministic function-fitting task.
To achieve more effective approximations of the solutions of PDEs, we design the DBR algorithm via the following manners: (i) the FBSDEs is discretized by the Euler scheme which is introduced by Zhang [36]; (ii) utilize the artificial neural networks equipped with the input variables $(t_{i},X_{i}^{\pi})$ to learn the conditional expectations at each time step $t_{i}$; (iii) adopt a step-by-step approach, optimizing the variables $Z_{i}$ first and then the variables $Y_{i}$ at each time step, rather than simultaneously optimizing the pairs $(Y_{i},Z_{i})$ ; (iv) introduce the loss functions (1.3) and (1.4) to train the deep feedforward neural networks.
@@
Moreover, the core of the DBR method is to approximate the conditional expectation through multi-path Monte Carlo averaging. This design shows some advantages when dealing with high-dimensional or complex-structured PDEs: (i) noise smoothing: multi-path averaging mitigates the impact of single-path Brownian noise $\Delta W_{i}$, reducing numerical fluctuations (e.g., the loss function of regression-based methods does not directly depend on $\Delta W_{i}$). (ii) generalization ability: the DBR method learns “local solutions of multi-paths” rather than “local solutions of single paths”, making them more likely to avoid overfitting. Especially, as the number of dimensions of PDEs with an unbounded solution and complex structure exceeds 10, the DBR method effectively mitigates the adverse effects of increasing dimensionality (see Tables 5 and 6). The method’s consistent performance at dimensions $d=15$ and $d=20$, particularly when contrasted with the breakdown of DBDP1, provides compelling evidence of its practical utility for high-dimensional problems.
@@
The remainder of this paper is organized as follows. In Section 2, some essential materials that are used in the whole paper are provided. The DBR method is designed in Subsection 3.1; The upper bound error of the proposed DBR method is rigorously derived theoretically in Subsection 3.2. In addition, we provide the convergence analysis for the proposed DBR method in Subsection 3.3. In Section 4, we extend the DBR scheme to variational inequalities and prove its corresponding convergence. We report the numerical results in Section 5. In the end, we compile the conclusion of this paper in Section 6.
@@ 2 Preliminaries
In this section, we present some materials that will be utilized to design the deep probabilistic numerical approximation method for the high-dimensional nonlinear PDEs, such as the probabilistic representation of the solutions to the PDEs, the function approximations in the class of neural networks and the Euler scheme for FBSDEs.
@@ 2.1 Deep neural networks as function approximators
2.2 Probabilistic representation for the solutions of PDEs
Our deep probabilistic numerical method for the PDEs (1.1) relies on FBSDEs representation of its solutions. The well-known nonlinear Feynman-Kac formula (see [31, 33, 32]) shows that on the filtered complete probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{P})$, the solution $(Y,Z)\in\mathbb{R}^{n}\times\mathbb{R}^{n\times d}$ of the FBSDEs
@@ 2.2 Probabilistic representation for the solutions of PDEs
related to the solution $u$ of the PDEs (1.1) via
$Y_{t}=u(t,X_{t}),\qquad Z_{t}=\sigma^{\top}(t,X_{t})D_{x}u(t,X_{t}),$ (2.2)
where $u(s,x)\in C_{b}^{1,2}([0,T]\times\mathbb{R}^{d})$; $(\mathcal{F}_{t})_{0\leq t\leq T}$ denotes the standard Brownian filtration; $W_{t}$ denotes a $d$-dimensional Brownian motion; $x_{0}\in\mathbb{R}^{d}$ represents a given initial value of the stochastic differential equation (SDE) in (2.1).
For readers’ convenience, before giving the time-discretization scheme of FBSDEs (2.1), we introduce the following notations. We denote the grid of the time interval $[0,T]$ by $\pi$, namely $\pi=\{0=t_{0}<t_{1}<\cdots<t_{N}=T\}$, where $t_{i}=ih,h=\frac{T}{N}$ for $i=0,1,\cdots,N,N\in\mathbb{N}^{+}$ and $\Delta W_{i}=W_{t_{i+1}}-W_{t_{i}}$ the $(i+1)$-th Brownian motion increment. Then the Euler time-discretization of the the backward stochastic differential equation (BSDE) in (2.1), at the mesh points $t_{i}$, is as follows
$Y_{i}^{\pi}=Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},Y_{i}^{\pi},Z_{i}^{\pi})-Z_{i}^{\pi}\Delta W_{i},$ (2.3)
@@
$\left\{\begin{aligned} Y_{i}^{\pi}=&\mathbb{E}_{i}\Big[Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},Y_{i}^{\pi},Z_{i}^{\pi})\Big],\\ Z_{i}^{\pi}=&\mathbb{E}_{i}\big[Y_{i+1}^{\pi}\frac{\Delta W_{i}^{\top}}{h}\big],\end{aligned}\right$. (2.4)

### 21 May 2026 v3

@@
Abstract
We propose a deep backward regression-based (DBR) scheme for solving high-dimensional nonlinear parabolic partial differential equations. Building on the DBDP method of Huré, Pham, and Warin [29], the proposed method reformulates the local backward losses through conditional expectations and trains the resulting regression problems sequentially in time. This conditional-expectation formulation replaces pathwise Brownian fluctuations in the Euler residual by their averaged effect and therefore provides an intrinsic variance-reduction mechanism before loss evaluation. In practice, the conditional expectations are approximated by local multi-path Monte Carlo averages, which leads to smoother training targets and improved numerical stability. Numerical experiments show that DBR performs competitively on standard high-dimensional benchmarks and is more stable than DBDP1 on the challenging unbounded benchmark considered in Example 2. Under an idealized population-loss minimization setting, we provide an error analysis and establish a half-order convergence result under suitable approximation and integrability assumptions. We also discuss an extension to variational inequalities.
keywords:
high dimensional nonlinear PDEs, backward stochastic differential equations, conditional expectations, deep neural networks, numerical approximations.
@@ 1 Introduction
In this paper, we develop a probabilistic deep learning scheme for approximating solutions of high-dimensional quasilinear parabolic PDEs of the form
@@
The pioneer works of Pardoux and Peng [33] and Peng [34] show that the quasilinear parabolic PDEs (1) are associated with Markovian forward backward stochastic differential equations (FBSDEs) due to the nonlinear Feynman-Kac formula. Since then, significant interest has been directed toward the probabilistic interpretation of PDE solutions. By these interpretations, researchers have developed probabilistic numerical methods to approximate PDE solutions (see [9, 10, 12] and references therein). Concurrently, the design of algorithms to solve FBSDEs has been intensively studied. A significant milestone in this classical numerical analysis line was established by Zhang in his seminal work [37]. This work provided a robust theoretical foundation. In fact, our framework is fundamentally rooted in the classical backward stochastic differential equations (BSDEs) numerical analysis, drawing particular inspiration from Zhang [37], which proposed the idea of solving BSDEs through backward induction on conditional expectations and proved the rate of convergence in the strong $L^{2}$ sense. The DBR method effectively extends these fundamental principles from traditional numerical method into the deep learning paradigm, aiming to handle the curse of dimensionality. Up to now, the proposed numerical algorithms to solve nonlinear FBSDEs have shown their remarkable performance with respect to the high-order problems. For instance, the Euler schemes with convergence rates of $\frac{1}{2}$ or $1$ are presented in [2, 20, 38]. Multi-step schemes [25, 26, 22, 39] achieve convergence rates greater than or equal to $1$, although they require terminal conditions over multiple steps. Furthermore, one-step schemes with convergence rates of at least $2$ have been proposed in [8, 36, 27].
@@
However, most of these numerical methods are unsuitable for solving nonlinear PDEs in dimensions greater than 4 (see [4]). To address this issue, some techniques such as the parallel computing method, the sparse grid method, and the variance reduction method (including the control variate methods, the importance sampling methods, and the multilevel Monte Carlo methods et al.) have proven effective for nonlinear PDEs with the dimension below 20 (see [21]). The high dimensional problem (that is the so-called curse of dimensionality) remains a major challenge in the numerical solutions of PDEs, even in the broader field of scientific computing.
@@
A major development occurred in 2017, when E, Han, and Jentzen introduced the deep BSDE method, one of the first deep learning-based numerical methods for high-dimensional nonlinear PDEs and BSDEs. Since then, many neural-network-based numerical methods have been proposed. They are designed by adjusting, reformulating or extending the deep BSDE method to solve high dimensional PDEs (see [13, 17, 29, 24, 30, 35, 15, 16] and many others) and these deep numerical algorithms have also shown their remarkable performance on the high-dimensional problems. Recently, Cai, Fang and Zhou [5, 7] introduce the SOC-MartNet method to solve ultra-high-dimensional quasilinear parabolic PDEs and it demonstrates the strong performance on large-scale benchmarks. Furthermore, they propose the deep random difference method in [6] to reduce variance and improve stability of the SOC-MartNet method.
@@
At the forefront of deep learning methodologies for high-dimensional PDEs, the seminal work by Hure, Pham, and Warin [29] proposed the DBDP1 method, which has become a leading paradigm in the current literature. It estimates the solutions and their gradient by the loss functions, for $i=N-1,N-2,\cdots,0$
$J_{i}^{HPW}(\theta_{i})=\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{Y}_{i+1}^{*}-hf(t_{i},X_{i}^{\pi},\mathcal{Y}_{i},\mathcal{Z}_{i})+\mathcal{Z}_{i}\Delta W_{i}\right|^{2}\right],$ (2)
@@
at each time step $t_{i}$, where $\mathcal{Y}_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ and $\mathcal{Z}_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n\times d}$ denote the sequences of network functions. Then update $(\mathcal{Y}_{i}^{*},\mathcal{Z}_{i}^{*})$ as the solutions of the local minimization problems. Thus, we only need to consider the local gradient as updating the parameter of neural networks, which significantly reduces the difficulty of backpropagation and solves the training difficulties brought by long-sequence time dependencies. The loss function (2) is based on the Euler discretization of the BSDE. It measures the residual of the Euler scheme which implies that during the training process, one can directly assess the accuracy of the scheme by observing whether the loss value approaches zero. But the random noise term $\mathcal{Z}_{i}\Delta W_{i}$ in (2) may lead to fluctuations especially when $\mathcal{Z}_{i}$ is large (such as high volatility or complex hedging requirements). Naturally, it may cause oscillations in the gradients which lead to divergence. Besides, the loss function $J_{i}^{HPW}(\theta_{i})$ is prone to overfitting the noise of the simulated paths, leading to poor generalization ability.
@@
$F_{y,i}(\theta_{y,i})=\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}^{*}+hf(t_{i},X_{i}^{\pi},\mathcal{Y}_{i},\mathcal{Z}_{i}^{*})\right]\right|^{2}\right],$ (3)
@@
$F_{z,i}(\theta_{z,i})=\mathbb{E}\left[\left|\mathcal{Z}_{i}-\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}^{*}\frac{\Delta W_{i}^{\top}}{h}\right]\right|^{2}\right],$ (4)
@@
where $\mathbb{E}_{i}\left[\cdot\right]=\mathbb{E}\left[\cdot\big|\mathcal{F}_{t_{i}}\right]$ is the conditional expectation. The above regression-based loss functions which utilize conditional expectations are different from those proposed in Remark 2.3 of Germain, Pham and Warin [17]. Unlike the DBDP method, which operates directly on the Euler-discretized residual susceptible to pathwise noise $\Delta W_{i}$, our DBR method exploits the conditional expectation representation central to classical frameworks like Zhang [37]. By replacing pathwise noise with its averaged effect, the DBR scheme achieves intrinsic variance reduction. This design choice aligns conceptually with the $L^{2}$-regularity established by Zhang [37], which provides the theoretical foundation for smoothing the martingale integrand $Z$ via step-process approximations. The accuracy of the DBR scheme can also be tested when computing at each time step the infimum of loss function, which should be equal to zero as the DBDP scheme in [29]. By explicitly incorporating the conditional expectations, the DBR scheme effectively performs a “denoising” step prior to loss evaluation. In terms of implementation, this is typically realized through multi-path Monte Carlo averaging. Its essence lies in replacing the simulated backward stochastic difference equations with their conditional expectation forms and transforming a projection-based stochastic optimization problem into a smoother deterministic function-fitting task.
The DBR algorithm is built on the following ingredients: (i) the FBSDE is discretized by an Euler-type scheme inspired by Zhang [37]; (ii) neural networks with inputs $(t_{i},X_{i}^{\pi})$ are used to approximate the conditional expectations at each time step; (iii) the $Z$-component is trained first and then fixed when training the $Y$-component; (iv) the conditional-expectation-based losses (3) and (4) are used for network training.
@@
Moreover, the core of the DBR method is to approximate the conditional expectation through multi-path Monte Carlo averaging. This design shows some advantages when dealing with high-dimensional or complex-structured PDEs: (i) noise smoothing: multi-path averaging mitigates the impact of single-path Brownian noise $\Delta W_{i}$, reducing numerical fluctuations (e.g., the loss function of regression-based methods does not directly depend on $\Delta W_{i}$). (ii) generalization ability: the DBR method learns local solutions of multi-paths rather than local solutions of single paths, making them more likely to avoid overfitting. In particular, for the unbounded benchmark with a complex structure considered in Example 2, the DBR method mitigates the adverse effects of increasing dimensionality more effectively than DBDP1 when $d=15$ and $d=20$; see Tables 5.5 and 5.6. The method’s consistent performance at these two dimensions, compared with the substantial deterioration of DBDP1 on the same benchmark, provides evidence of its practical utility for this class of challenging high-dimensional problems.
@@
The remainder of this paper is organized as follows. In Section 2, some essential materials that are used in the whole paper are provided. The DBR method is designed in Subsection 3.1. An error analysis for the proposed DBR method is presented in Subsection 3.2 under the population-loss convention. In addition, under suitable approximation and integrability assumptions, we establish a half-order convergence result for the proposed DBR method in Subsection 3.3. In Section 4, we extend the DBR scheme to variational inequalities and prove its corresponding convergence. We report the numerical results in Section 5. In the end, we compile the conclusion of this paper in Section 6.
@@ 2 Preliminaries
This section collects the notation and preliminary results used throughout the paper, including the probabilistic representation of PDE solutions, neural-network function classes, and the Euler discretization of FBSDEs.
@@ 2.1 Deep neural networks as function approximators
2.2 Probabilistic representation for the PDE solutions
Our deep probabilistic numerical method for the PDEs (1) relies on FBSDEs representation of its solutions. The well-known nonlinear Feynman-Kac formula (see [32, 34, 33]) shows that on the filtered complete probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{P})$, the solution $(Y,Z)\in\mathbb{R}^{n}\times\mathbb{R}^{n\times d}$ of the FBSDEs
@@ 2.2 Probabilistic representation for the PDE solutions
related to the solution $u$ of the PDEs (1) via
$Y_{t}=u(t,X_{t}),\qquad Z_{t}=\sigma^{\top}(t,X_{t})D_{x}u(t,X_{t}),$ (6)
where $u$ is a sufficiently smooth classical solution. In order to include unbounded terminal functions and unbounded solutions, we shall use the following polynomial growth class instead of the bounded class $C_{b}^{1,2}$. We write
$u\in C_{p}^{1,2}([0,T]\times\mathbb{R}^{d})$
if $u\in C^{1,2}([0,T]\times\mathbb{R}^{d})$ and there exist constants $C_{u}>0$ and $q_{u}\geq 0$ such that
$|u(t,x)|+|D_{x}u(t,x)|+|D_{x}^{2}u(t,x)|\leq C_{u}\left(1+|x|^{q_{u}}\right),\qquad(t,x)\in[0,T]\times\mathbb{R}^{d}.$ (7)
The bounded case $C_{b}^{1,2}$ corresponds to $q_{u}=0$. Under this regularity and the integrability assumptions stated below, the nonlinear Feynman-Kac formula yields
$Y_{t}=u(t,X_{t}),\qquad Z_{t}=\sigma^{\top}(t,X_{t})D_{x}u(t,X_{t}).$
Here $(\mathcal{F}_{t})_{0\leq t\leq T}$ denotes the standard Brownian filtration; $W_{t}$ denotes a $d$-dimensional Brownian motion; $x_{0}\in\mathbb{R}^{d}$ represents a given initial value of the stochastic differential equation (SDE) in (5).
For readers’ convenience, before giving the time-discretization scheme of FBSDEs (5), we introduce the following notations. We denote the grid of the time interval $[0,T]$ by $\pi$, namely $\pi=\{0=t_{0}<t_{1}<\cdots<t_{N}=T\}$, where $t_{i}=ih,h=\frac{T}{N}$ for $i=0,1,\cdots,N,N\in\mathbb{N}^{+}$ and $\Delta W_{i}=W_{t_{i+1}}-W_{t_{i}}$ the $(i+1)$-th Brownian motion increment. Then the Euler time-discretization of the backward stochastic differential equation (BSDE) in (5), at the mesh points $t_{i}$, is as follows
$Y_{i}^{\pi}=Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},Y_{i}^{\pi},Z_{i}^{\pi})-Z_{i}^{\pi}\Delta W_{i},$ (8)
@@
$\left\{\begin{aligned} Y_{i}^{\pi}=&\mathbb{E}_{i}\Big[Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},Y_{i}^{\pi},Z_{i}^{\pi})\Big],\\ Z_{i}^{\pi}=&\mathbb{E}_{i}\big[Y_{i+1}^{\pi}\frac{\Delta W_{i}^{\top}}{h}\big],\end{aligned}\right$. (9)
@@
Since the first equation in the above conditional expectation formula is implicit with respect to $Y_{i}^{\pi}$, we impose a standard small time-step condition to ensure its well-posedness. Let $L_{y}$ denote a Lipschitz constant of $f$ with respect to the $y$-variable. Under Assumption (iii) below, one can take $L_{y}\leq L$. We assume throughout that
$hL_{y}<1.$ (10)
Indeed, for fixed $i$, $X_{i}^{\pi}$, and $Z_{i}^{\pi}$, the map
$y\longmapsto\mathbb{E}_{i}\left[Y_{i+1}^{\pi}\right]+hf(t_{i},X_{i}^{\pi},y,Z_{i}^{\pi})$
is a contraction on $\mathbb{R}^{n}$ whenever (10) holds. Hence the implicit equation for $Y_{i}^{\pi}$ admits a unique solution at each time step.